# Serialized Speech Information Guidance with Overlapped Encoding Separation for Multi-Speaker Automatic Speech Recognition

### v1

@@ 1 Introduction
Automatic speech recognition gets impressive performance with the development of deep learning [1, 2, 3, 4]. The word error rate (WER) of ASR for single speaker conditions has achieved the level of human transcribers [4, 5], even when faced with many complex scenarios, such as noise [6, 7, 8, 9, 10, 11]. However, compared to the additive noise and reverberation, the inference from other speakers, known as the cocktail problem, has a more severe effect on ASR [12, 13, 14]. As a result, ASR performance has dramatic degradation under multi-speaker scenarios [13].
Recently, extensive research has been conducted on multi-speaker ASR [13, 14, 15, 16]. It is intuitive to decompose the multi-speaker ASR task to speech separation and then recognition [17, 18]. The pipeline consists of a speech separation front-end [19, 7, 20] followed by a recognizer [17]. However, speech separation front-ends, especially the single-channel methods, often bring speech information loss and distortion issues, which harms ASR [17, 21]. With the development of end-to-end ASR, the ASR back-end already has some abilities to handle multi-speaker conditions [22, 23]. Utterance-level permutation invariant training (uPIT) [22] is popular for multi-speaker ASR [24]. During training, all possible permutations of speakers are needed to compute loss, and the smallest loss is used for backpropagation [24]. In the uPIT-based ASR, the number of output layers constrains the maximum number of speakers [25], and training becomes computationally complex with more speakers [25].
@@ 2.3 Self-Supervised Learning-based Feature Extraction
Feature extraction with self-supervised learning (SSL) shows powerful performance improvement for ASR [8]. The outputs of the SSL module replace the traditional features like Mel-Frequency Cepstral Coefficients (MFCC) and Filter Bank (F-Bank). The SSL-based feature extraction module typically consists of convolution layers and transformer, e.g., WavLM [32], HuBERT [33]. With massive pretraining data, the SSL models already contained the capability to convert the speech waveform into the acoustic hidden units. Moreover, some SSL models [32, 33] demonstrate robustness against noise or other inference speakers by incorporating noisy simulation. The other parts of the ASR system are the same as the other E2E ASR systems.
@@ 4.1 Datasets and Experimental Settings
We used the LibriMix dataset [35] to evaluate the model performance. It used the train-clean-100, train-clean-360, dev-clean, and test-clean subsets from the LibriSpeech dataset [36] as the clean speech. The noise samples were taken from WHAM! dataset [37]. We used the official scripts [^1] https://github.com/JorisCos/LibriMix to synthesize Libri2Mix and Libri3Mix. We used the offset file to make different speaking start times for multiple speakers. The two-speaker offset files followed the official ESPnet setting , and the three-speaker offset files were created by ourself [^2] https://github.com/anonymous4blind/LibriMix-repo.

### v3

@@ 1 Introduction
Automatic speech recognition gets impressive performance with the development of deep learning [1, 2, 3, 4, 5, 6]. The word error rate (WER) of ASR for single speaker conditions has achieved the level of human transcribers [3, 7], even when faced with many complex scenarios, such as noise [8, 9, 10, 11, 12]. However, compared to the additive noise and reverberation, the inference from other speakers, known as the cocktail problem, has a more severe effect on ASR [13, 14, 15]. As a result, ASR performance has dramatic degradation under multi-speaker scenarios [14].
Recently, extensive research has been conducted on multi-speaker ASR [14, 15]. It is intuitive to decompose the multi-speaker ASR task to speech separation and then recognition [16, 17]. The pipeline consists of a speech separation front-end [18, 19] followed by a recognizer [16]. However, speech separation front-ends, especially the single-channel methods, often bring speech information loss and distortion issues, which harms ASR [16, 20]. With the development of end-to-end ASR, the ASR back-end already has some abilities to handle multi-speaker conditions [21, 22]. Utterance-level permutation invariant training (uPIT) [21] is popular for multi-speaker ASR [23]. During training, all possible permutations of speakers are needed to compute loss, and the smallest loss is used for backpropagation [23]. In the uPIT-based ASR, the number of output layers constrains the maximum number of speakers [24], and training becomes computationally complex with more speakers [24].
@@ 2.3 Self-Supervised Learning-based Feature Extraction
Feature extraction with self-supervised learning (SSL) shows powerful performance improvement for ASR [9]. The outputs of the SSL module replace the traditional features like Mel-Frequency Cepstral Coefficients (MFCC) and Filter Bank (F-Bank). The SSL-based feature extraction module [31, 32, 33] typically consists of convolution layers and transformer, e.g., WavLM [34], HuBERT [35]. With massive pretraining data, the SSL models already contained the capability to convert the speech waveform into the acoustic hidden units. Moreover, some SSL models [34, 35] demonstrate robustness against noise or other inference speakers by incorporating noisy simulation. The other parts of the ASR system are the same as the other E2E ASR systems.
@@ 4.1 Datasets and Experimental Settings
We used the LibriMix dataset [37] to evaluate the model performance. It used the train-clean-100, train-clean-360, dev-clean, and test-clean subsets from the LibriSpeech dataset [38] as the clean speech. The noise samples were taken from WHAM! dataset [39]. We used the official scripts [^1] https://github.com/JorisCos/LibriMix to synthesize Libri2Mix and Libri3Mix. We used the offset file to make different speaking start times for multiple speakers. The two-speaker offset files followed the official ESPnet setting  [^2] https://github.com/espnet/espnet/tree/master/egs2/librimix/sot_asr1 , and the three-speaker offset files were created by ourself.